OpenStudy (anonymous):

i need to learn how to find a equation of the line with the given y-intercept and slop m -4.5,m=2.5

OpenStudy (jmark):

The general equation of the line having slope m is y = mx + c Given y intercept(means the value of c) = -4.5 and slope (m) = 2.5 (Note: to find the y-intercept we put x = 0, so y = 0 +c or y = c) By substituting given points in the above eqaution, we get y = 2.5x - 4.5 (which is required eqaution) CBSE 11th Commerce Syllabus - http://scoremore.beep.com/cbse-11th-commerce-syllabus-2014-07-31-1.htm